### v5

@@ 1 Introduction
The other extreme is when just enough cycles are present for one atom to be flexible. This case was handled by Alm, Maddux & Manske [AMM08], who exhibited a representation of the algebra $A_{n}$ obtained from splitting the non-flexible diversity atom of $6_{7}$ into $n$ symmetric atoms. In particular, this construction yielded a representation of $A_{2}=32_{65}$ over a set of size $416,714,805,91$ (here, we use Maddux’s [Mad06] numbering for relation algebras such as $6_{7}$ and $32_{65}$). Dodd & Hirsch [DH13] subsequently improved the upper bound of the minimum representation size of $32_{65}$ to $63,432,274,896$. This was subsequently improved to 8192 by J.F. Alm & D. Sexton (unpublished), and later $3432$ by [AA19]. Finally, in [ALM+22], the authors exhibited a representation over a set of size $1024$ for $32_{65}$, as well as the first polynomial upper bound on $\min(\operatorname{Spec}(A_{n}))$ (where $\operatorname{Spec}(A_{n})$ is the spectrum taken over all square representations).
@@ 6.2 $33_{65}$
Consider the red edge $x_{2}x_{5}$, which has $9$ needs. The b-g need is met by $x_{0}$, and the g-g need is met by $x_{1}$. Now the r-r need is met by $x_{3},x_{4},x_{6},x_{7},x_{10}$. We can meet the r-g need with $x_{8}$ and an arbitrary need with $x_{9}$. This leaves $4$ unsatisfied needs, necessitating $4$ additional points. So we require at least $15$ points. ∎
@@
If $16\leq n\leq 23$, then $n\not\in\operatorname{Spec}(33_{65})$.
@@ 6.3 $34_{65}$
We again phrase our results in the language of graph theory (see Remark 2.12). We use the color red to correspond to the flexible atom, green to correspond to the atom $b$, and blue to correspond to the atom $c$. Thus, the forbidden triangles are the blue-blue-green and blue-green-green ones.
@@ 8 Conclusion
Conjecture 8.1 (Strong Flexible Atom Conjecture).
Every finite integral RA with a flexible atom is representable over a Comer scheme.

### v6

@@ 1 Introduction
The other extreme is when just enough cycles are present for one atom to be flexible. This case was handled by Alm, Maddux & Manske [AMM08], who exhibited a representation of the algebra $A_{n}$ obtained from splitting the non-flexible diversity atom of $6_{7}$ into $n$ symmetric atoms. [^2] Here, we use Maddux’s [Mad06] numbering for relation algebras such as $6_{7}$ and $32_{65}$. Note that $6_{7}$ is the relation algebra with two symmetric atoms $a,b$, where $a$ is flexible and the sole forbidden diversity cycle is $bbb$. We refer to Table 1 for $32_{65}$. In particular, this construction yielded a representation of $A_{2}=32_{65}$ over a set of size $416,714,805,91$. Dodd & Hirsch [DH13] subsequently improved the upper bound of the minimum representation size of $32_{65}$ to $63,432,274,896$. This was subsequently improved to 8192 by J.F. Alm & D. Sexton (unpublished), and later $3432$ by [AA19]. Finally, in [ALM+22], the authors exhibited a representation over a set of size $1024$ for $32_{65}$, as well as the first polynomial upper bound on $\min(\operatorname{Spec}(A_{n}))$ (where $\operatorname{Spec}(A_{n})$ is the spectrum taken over all square representations).
@@ 6.2 $33_{65}$
Consider the red edge $x_{2}x_{5}$, which has $9$ needs. The g-b need is met by $x_{0}$, and the g-g need is met by $x_{1}$. Now the r-r need is met by $x_{3},x_{4},x_{7},x_{8},x_{10}$. We can meet the r-g need with $x_{6}$ and an arbitrary need with $x_{9}$. This leaves $4$ unsatisfied needs, necessitating $4$ additional points. So we require at least $15$ points. ∎
@@
If $15\leq n\leq 23$, then $n\not\in\operatorname{Spec}(33_{65})$.
@@ 6.3 $34_{65}$
We again phrase our results in the language of graph theory (see Remark 2.12). We use the color red to correspond to the flexible atom, blue to correspond to the atom $b$, and green to correspond to the atom $c$. Thus, the forbidden triangles are the blue-blue-green and blue-green-green ones.
@@ 8 Conclusion
Conjecture 8.1 (Strong Commutative Flexible Atom Conjecture).
Every finite commutative integral RA with a flexible atom is representable over a Comer scheme.
Remark 8.2.
The assumption that the relation algebras are commutative is indeed necessary, and we discovered this only after the conference version [AAL23] where we omitted the commutative hypothesis. The algebra $1284_{1316}$ has diversity atoms $a$, $b$, $r$, and $\breve{r}$, where $a$ is flexible and the only forbidden diversity cycle is $rbr$. However, $1284_{1316}$ is not commutative, and therefore not representable over an Abelian group. In particular, Comer’s method does not yield a group representation.
It would even be of interest to look at the special case of symmetric relation algebras. Note that if all our diversity atoms are symmetric, then the relation algebra is commutative. However, the first and third authors had considerable success leveraging Comer’s method to provide finite representations for non-symmetric, but commutative relation algebras [AL23].